# The Baryonic Tully-Fisher Relation of HI-bearing Low Surface Brightness Galaxies Implies Their Formation Mechanism

### v2

@@ 1 Introduction
Low surface brightness galaxies (LSBGs) constitute a special population of galaxies with central surface brightness at least one magnitude fainter than that of the sky background (i.e., $B$-band surface brightness $\rm\geq 22.5\ mag\cdot arcsec^{-2}$; Impey & Bothun, 1997; Bothun et al., 1997). In the fields, LSBGs have been observed to exhibit higher HI fractions (Du et al., 2015; He et al., 2020), lower star formation rates (Wyder et al., 2009; Rong et al., 2020b), and low metallicities (Kuzio de Naray et al., 2004). Also, LSBGs tend to appear in low-density environments (Pérez-Montaño & Cervantes Sodi, 2019; Mo et al., 1994; Galaz et al., 2011).
@@
The formation and evolution mechanisms of LSBGs, which play a crucial role in refining our current models of galaxy formation and evolution, have gathered much attention ever since their initial detection. Early investigations often regarded LSBGs as galaxies formed in low-density dark matter halos (Dekel & Silk, 1986; McGaugh, 1992; Mo et al., 1994). However, more recent studies have presented more alternative scenarios. One of the most classical theories of the formation mechanism of LSBGs is that they are galaxies hosted in dark matter halos belonging to the high-spin tail of the spin distribution. In theoretical works and simulations, the halos of LSBGs are found to have higher spin parameters compared with high surface brightness galaxies, regardless of whether they are relatively massive (Mo et al., 1998; Kim & Lee, 2013; Kulier et al., 2020; Pérez-Montaño et al., 2022), or belong to ultra-diffuse galaxies (UDGs, Rong et al., 2017; Amorisco & Loeb, 2016; Rong et al., 2024) – a specific subset of LSBGs of smaller masses. Additionally, some studies suggest that low-mass LSBGs may represent failed $L^{\rm*}$ galaxies (van Dokkum et al., 2015a, b), or originate from early co-planar mergers (Wright et al., 2021) or stellar feedback processes (Chan et al., 2018; Di Cintio et al., 2019), or birth during tidal interactions (Rong et al., 2020a; Carleton et al., 2019), or form through multiple pathways (Papastergis et al., 2017). High-mass LSBGs, alternatively, may form through mergers (Saburova et al., 2018; Di Cintio et al., 2019; Zhu et al., 2023), two-stage process with external gas accretions (Saburova et al., 2021), or dynamical evolution driven by bars (Noguchi, 2001).
@@
Nevertheless, despite extensive efforts, our present comprehension of the formation and evolution of LSBGs is still limited. One of the key reasons is that studies on the mass distributions and baryonic fractions in LSBGs have not achieved a consensus, and a comprehensive scenario has not been established to constrain the current formation and evolution models of LSBGs. Many studies have shown that the dynamics of LSBGs may be dominated by dark matter (e.g. de Blok & McGaugh, 1997; Mowla et al., 2017), yet UDGs may exhibit a deficiency in dark matter content (e.g. van Dokkum et al., 2018; Mancera Piña et al., 2019). As for the stellar-to-halo mass relation ($M_{\rm*}-M_{\rm halo}$), while Prole et al. (2019) propose that LSBGs form a continuous extension of typical dwarfs in the $M_{\rm*}-M_{\rm halo}$ diagram, other studies reveal significant deviations for UDGs from the typical $M_{\rm*}-M_{\rm halo}$ relation (e.g. van Dokkum et al., 2018). Additionally, some giant LSBGs appear to be dominated by baryons at their central regions, akin to the high surface brightness galaxy (HSBG) counterparts (Lelli et al., 2010); the stellar-to-halo mass relation of the giant LSBGs and HSBGs also exhibits a similarity. All of these confusing signs of diversity in matter distribution among LSBGs could offer valuable insights into their varied formation and evolution pathways. Consequently, a comprehensive understanding of the matter distributions in LSBGs is essential for refining our current models of galaxy formation and evolution.
@@
The baryonic Tully-Fisher relation (BTFR) is a correlation between the rotational velocity, $v_{\rm rot}$, and the baryonic mass, $M_{\rm b}$, of galaxies. Over the past two decades, many studies have shown that late-type galaxies and dwarf galaxies consistently adhere to a power-law form of BTFR, despite their wide ranges of baryonic mass, luminosity, and size. BTFR could be written as,
$M_{\rm b}\propto v_{\rm rot}^{\beta},$ (1)
where the logarithmic slope, $\beta$, typically falls within a range from 3 to 4 (e.g. McGaugh et al., 2000; Begum et al., 2008; Papastergis et al., 2016; Lelli et al., 2016b; Sales et al., 2017; Goddy et al., 2021). Notably, some studies have also discovered that early-type galaxies may also follow the BTFR of late-type galaxies (den Heijer et al., 2015; Serra et al., 2012).
@@
In the context of $\rm\Lambda CDM$ paradigm, the presence of BTFR is considered as a correlation between the virial velocity of the dark matter halo and the baryonic mass enclosed within that halo (McGaugh, 2012; Mo et al., 1998). Thus, it can serve as a valuable tool to probe the distribution of baryonic matter and dark matter within galaxies. Furthermore, simulations have highlighted the significance of feedback processes in shaping the observed BTFR (e.g. Governato et al., 2010; Piontek & Steinmetz, 2011; Dutton, 2012), suggesting that investigating the BTFR can serve as a valuable means to constrain the feedback mechanisms at work within galaxies. Additionally, BTFR is highly related to other dynamical relationships such as the radial acceleration relation (e.g. McGaugh et al., 2016; Lelli et al., 2017). Exploring the BTFR could also provide insights into new gravitational theories, as a BTFR with $\beta=4$ is well predicted by the MOdified Newtonian Dynamics (MOND, Milgrom, 1983). Given its multifaceted implications, the BTFR remains one of the most extensively studied dynamical relations in astrophysics.
@@
As anticipated, recent studies examining the BTFR of LSBGs also show some potential signs of tension. The previous studies of, for instance, Lelli et al. (2016b) (see also Zwaan et al., 1995), have explored that the residuals of BTFR do not change with galactic surface brightness, while some of the recent investigations (Hu et al., 2023; Karunakaran et al., 2020; Rong et al., 2024) reveal that UDGs may significantly deviate from the BTFR of typical late-type galaxies. Therefore, a thorough and meticulous examination of the BTFR for LSBGs becomes essential. Such an analysis may not only indicate whether the matter distribution within LSBGs aligns with that of typical late-type and dwarf galaxies, but also provide valuable constraints on the formation mechanisms of LSBGs.
@@
In this work, we will first introduce our sample selection in Section 2, and then describe our methods for calculating $M_{\rm b}$ and $v_{\rm rot}$ in Section 3. We will then present our results in Section 4, and finally discuss the potential selection biases, the radial acceleration relation, and the possible formation mechanism of LSBGs in Section 5. The summary is in Section 6.
@@ 2 Sample Selection
From the parent sample, we first remove LSBGs with $g$-band apparent axis ratios $b/a>0.72$, approximately corresponding to the threshold of inclination angles of $i\lesssim 45^{\circ}$, because that the face-on galaxies may introduce large uncertainties in the measurements of the rotation velocities (see equation (10)). We then exclude LSBGs with low $\rm HI$ spectral signal-to-noise ratios ${\rm SNR}<10$, and visually remove LSBGs with suspicious HI spectra.
Due to the bad spatial resolution of ALFALFA (Durbala et al., 2020, being $3.8^{\prime}\times 3.3^{\prime}$;), if a galaxy has a close companion galaxy, its $\rm HI$ spectrum may be contaminated by the companion. To eliminate this possibility, we remove the sample galaxies with neighbouring galaxies within the projected radii of $3.8^{\prime}$ (i.e., approximately the beam size of Arecibo) and radial velocity differences of $\Delta v<500{\rm km/s}$. The redshift of a LSBG is estimated from the central velocity of its $\rm HI$ spectrum, while the redshift of the neighbouring galaxies comes from the SpecObj and Photoz databases of SDSS. Finally, 204 LSBGs are left in our sample.
We also calculate the kurtosis coefficients, $k_{\rm 4}$, of the HI spectra (Papastergis et al., 2016; El-Badry et al., 2018) for our sample galaxies, as
@@
In equations (3) and (4), $\nu$ and $F(\nu)$ represent the frequency and the flux density, respectively, and $\bar{\nu}$ is defined as,
@@
Here, $\nu_{\rm min}$ and $\nu_{\rm max}$ correspond to the minimal and maximum frequencies between which $W_{20}$ (see Section 3 for the definition of $W_{20}$) is measured, respectively. The coefficient $k_{\rm 4}$ quantifies the “peakiness” of a spectrum and thus can indicate whether a spectrum is single-peaked or double-horned. The single-peaked galaxies may have kinematics strongly influenced by velocity dispersion (El-Badry et al., 2018), or HI disks that are not extended enough to trace the flat part of the rotation curve (Papastergis et al., 2016). As a consequence, their HI-widths might not represent their rotation velocities (McGaugh, 2012; Verheijen, 1997). Therefore, we perform a cut at the HI spectrum, $k_{\rm 4}<-1$, in order to remove the single-peaked galaxies.
@@ 3 Method
We study the BTFRs of LSBGs and HSBGs by directly estimating the baryonic masses, $M_{\rm b}$, and rotation velocities, $v_{\rm rot}$, of the galaxies in the two samples.
The baryonic mass of a galaxy in our samples is mainly composed of the mass of stars and the mass of cold gas. The stellar mass $M_{\rm*}$ can be estimated with the galaxy luminosity in the $W1$-band (denoted as $L_{W1}$) from the Wide-field Infrared Survey Explorer (Wright et al., 2010), using the method proposed by McGaugh & Schombert (2015), as
@@
As for the mass of the cold gas, $M_{\rm g}$, we estimate it by adding up the masses of HI and helium,
$M_{\rm g}=1.33M_{\rm HI},$ (7)
where $M_{\rm HI}$ is the HI mass. Equation (7) will certainly underestimate the mass of the gas since it neglects the mass of the molecular gas and hot ionized gas. However, the molecular gas fraction in the total gas component is low and thus negligible (McGaugh, 2012). Also, the hot ionized gas usually resides outside the HI disk where our rotational velocity is measured (Dutton, 2012), so it can be neglected. Therefore, approximately we estimate the baryonic mass of a galaxy as,
$M_{\rm b}=M_{\rm*}+M_{\rm g}.$ (8)
For the rotation velocity, we estimate it using the $20\%$ peak width of a HI spectrum (Guo et al., 2020; Hu et al., 2023; Rong et al., 2024, $W_{\rm 20}$;), as,
@@
where $i$ is the inclination angle. $W_{20}$ employed in equation (9) has been corrected for instrumental broadening and redshift (Guo et al., 2020; Kent et al., 2008). In equation (9), we assume an alignment between the inclination angle of the stellar disk and HI disk in a galaxy, and then use the inclination angle of the stellar disk to estimate the rotation velocity. The inclination is estimated from the apparent axis ratio, as (e.g. Guo et al., 2020; Hu et al., 2023; Du et al., 2019; Rong et al., 2024; Begum et al., 2008),
$\sin i$ $=\sqrt{\frac{1-(b/a)^{2}}{1-q_{0}^{2}}},$ (10)
where $q_{0}$ denotes the intrinsic axis ratio seen edge-on. We take $q_{0}=0.2$, which is commonly used in the previous studies (e.g. Du et al., 2019; Guo et al., 2020; Giovanelli et al., 1997; Tully et al., 2009). For those edge-on galaxies with $g$-band $b/a\leq 0.2$, we set $\sin i=1$. Considering that the stellar disk and gas disk in galaxies may not be perfectly co-planar and often exhibit a small inclination difference of ${\rm\delta}i<20^{\circ}$ (e.g. Gault et al., 2021), we assume that the misalignment angle between the gas disk and the stellar disk, ${\rm\delta}i$, follows a Gaussian distribution centred at $0^{\circ}$ with a standard deviation of $\sigma=20^{\circ}$. We treat $\sigma=20^{\circ}$ as the uncertainty associated with $i$ in our study.
4 Results
In Figure 1, we observe that the distributions of the LSBGs and HSBGs exhibit no significant deviation in the $v_{\rm rot}-M_{\rm b}$ diagram. Subsequently, we impose a power-law equation,

### v3

@@ 1 Introduction
Low surface brightness galaxies (LSBGs) constitute a special population of galaxies with central surface brightness at least one magnitude fainter than that of the sky background (i.e., $B$-band surface brightness $\rm\geq 22.5\ mag\cdot arcsec^{-2}$; Impey & Bothun, 1997; Bothun et al., 1997). Many isolated LSBGs have been observed to exhibit higher HI fractions (e.g., Catinella et al., 2018; Schombert & McGaugh, 2021; Du et al., 2015; He et al., 2020). However, some studies (e.g., Martin et al., 2019; Galaz et al., 2002) alternatively found that some LSBGs exhibit the same or even smaller gas fractions compared to high surface brightness galaxies (HSBGs). LSBGs usually have low star formation rates (Wyder et al., 2009; Rong et al., 2020b), low metallicities (Kuzio de Naray et al., 2004; Schombert & McGaugh, 2021), and tend to appear in low-density environment (Pérez-Montaño & Cervantes Sodi, 2019; Mo et al., 1994; Galaz et al., 2011).
@@
The formation mechanism of LSBGs, which plays a crucial role in refining our empirical galaxy formation models, have gathered much attention ever since their initial detection. Early studies often regarded LSBGs as galaxies formed in low-density dark matter halos (Dekel & Silk, 1986; McGaugh, 1992; Mo et al., 1994). However, recent studies have presented more alternative scenarios. One classical formation scenario is that LSBGs are galaxies hosted in dark matter halos with higher halo spin. In cosmological simulations, the halos of LSBGs are found to have higher spin parameters compared with HSBGs, regardless of whether they are relatively massive (Mo et al., 1998; Kim & Lee, 2013; Kulier et al., 2020; Pérez-Montaño et al., 2022), or belong to ultra-diffuse galaxies (UDGs, Rong et al., 2017; Amorisco & Loeb, 2016; Rong et al., 2024) – a specific subset of LSBGs of smaller masses (e.g., van Dokkum et al., 2015a, 2018; Mancera Piña et al., 2018, 2019). Additionally, some studies suggest that low-mass LSBGs may represent failed $L^{\rm*}$ galaxies (van Dokkum et al., 2015a, b), or originate from early co-planar mergers (Wright et al., 2021) or stellar feedback processes (Chan et al., 2018; Di Cintio et al., 2019), or birth during tidal interactions (Rong et al., 2020a; Carleton et al., 2019), or form through multiple pathways (Papastergis et al., 2017). High-mass LSBGs, alternatively, may form through mergers (Saburova et al., 2018; Di Cintio et al., 2019; Zhu et al., 2023), two-stage process with external gas accretions (Saburova et al., 2021), or dynamical evolution driven by bars (Noguchi, 2001).
@@
Nevertheless, despite extensive efforts, our present comprehension of the formation and evolution of LSBGs is still limited. One of the key reasons is that studies on the mass distributions and baryonic fractions in LSBGs have not achieved a consensus, and a comprehensive scenario has not been established to constrain the current formation and evolution models of LSBGs. Many studies have shown that the dynamics of LSBGs may be dominated by dark matter (e.g., de Blok & McGaugh, 1997; Mowla et al., 2017; Swaters et al., 2003; Read et al., 2017; Mancera Piña et al., 2022), yet UDGs may exhibit a deficiency in dark matter content (e.g., van Dokkum et al., 2018; Mancera Piña et al., 2019). As for the stellar-to-halo mass relation ($M_{\rm*}-M_{\rm h}$), while Prole et al. (2019) propose that LSBGs form a continuous extension of typical dwarfs in the $M_{\rm*}-M_{\rm h}$ diagram, other studies reveal significant deviations for UDGs from the typical $M_{\rm*}-M_{\rm h}$ relation (e.g., van Dokkum et al., 2018). Additionally, some giant LSBGs appear to be dominated by baryons at their central regions, akin to the HSBG counterparts (e.g., Lelli et al., 2010; Saburova et al., 2021, 2019); the stellar-to-halo mass relation of the giant LSBGs and HSBGs also exhibits a similarity. All of these confusing signs of diversity in matter distribution among LSBGs could offer valuable insights into their varied formation and evolution pathways. Consequently, a comprehensive understanding of the matter distributions in LSBGs is essential for refining our current models of galaxy formation and evolution.
@@
The baryonic Tully-Fisher relation (BTFR) is a correlation between the rotational velocity, $v_{\rm rot}$, and the baryonic mass, $M_{\rm b}$, of galaxies. Over the past two decades, many studies have shown that late-type galaxies and dwarf galaxies consistently adhere to a power-law form of BTFR, despite their wide ranges of baryonic mass, luminosity, and size. BTFR could be written as, $M_{\rm b}\propto v_{\rm rot}^{\beta}$, where the logarithmic slope, $\beta$, typically falls within a range from 3 to 4 (e.g., McGaugh et al., 2000; Verheijen, 2001; Begum et al., 2008; Papastergis et al., 2016; Lelli et al., 2016a; Sales et al., 2017; Ponomareva et al., 2018; Goddy et al., 2023). Notably, some studies have also discovered that early-type galaxies may also follow the BTFR of late-type galaxies (den Heijer et al., 2015; Serra et al., 2012).
@@
In the context of $\rm\Lambda CDM$ paradigm, the presence of BTFR is considered as a correlation between the virial velocity of the dark matter halo and the baryonic mass enclosed within that halo (McGaugh, 2012; Mo et al., 1998). Thus, it can serve as a valuable tool to probe the distribution of baryonic matter and dark matter within galaxies. Furthermore, simulations have highlighted the significance of feedback processes in shaping the observed BTFR (e.g., Governato et al., 2010; Piontek & Steinmetz, 2011; Dutton, 2012), suggesting that investigating the BTFR can serve as a valuable means to constrain the feedback mechanisms at work within galaxies. Additionally, BTFR is highly related to other dynamical relationships such as the radial acceleration relation (e.g., McGaugh et al., 2016; Lelli et al., 2017). Exploring the BTFR could also provide insights into new gravitational theories, as a BTFR with $\beta=4$ is well predicted by the MOdified Newtonian Dynamics (MOND, Milgrom, 1983). Given its multifaceted implications, the BTFR remains one of the most extensively studied dynamical relations in astrophysics.
@@
As anticipated, recent studies examining the BTFR of LSBGs also show some potential signs of tension. Previous studies have found that the BTFR does not depend on galactic surface brightness or other galactic properties(e.g., Lelli et al., 2016a; Ponomareva et al., 2018; see also Zwaan et al., 1995). However, some recent investigations (e.g., Mancera Piña et al., 2019; Guo et al., 2020; Hu et al., 2023; Karunakaran et al., 2020; Rong et al., 2024) reveal that, the particular subsample of LSBGs with large disk sizes, i.e., UDGs, may significantly deviate from the BTFR of typical late-type galaxies, indicative of the possible dependence on size (Mancera Piña et al., 2020). Therefore, a thorough and meticulous examination of the BTFR for LSBGs becomes essential. Such an analysis may not only indicate whether the matter distribution within LSBGs aligns with that of typical late-type and dwarf galaxies, but also provide valuable constraints on the formation mechanisms of LSBGs.
@@
In this work, we will first introduce our sample selection in Section 2, and then describe our methods for calculating $M_{\rm b}$ and $v_{\rm rot}$ in Section 3. We will then present our results in Section 4, and finally discuss the potential selection biases and the possible formation mechanism of LSBGs in Section 5. The summary is in Section 6.
@@ 2 Sample Selection
$\mu_{0}(B)=\mu_{0}(g)+0.47\times[\mu_{0}(g)-\mu_{0}(r)]+0.17,$ (1)
to select LSBGs. The authors finally selected a sample of LSBGs with the $B$-band apparent central surface brightness of $22.5<\mu_{0}(B)<27.64\ {\rm mag/arcsec^{2}}$, which is treated as the parent sample and used in this work to study the BTFR of LSBGs.
From the parent sample, we first remove LSBGs with $g$-band apparent axis ratios $b/a>0.72$, approximately corresponding to the threshold of inclination angles of $i\lesssim 45^{\circ}$, because the face-on galaxies may introduce large uncertainties in the measurements of the rotation velocities (see equation (7)). We then exclude LSBGs with low $\rm HI$ spectral signal-to-noise ratios ${\rm SNR}<10$ (see Haynes et al., 2018 for the definition of SNR). Furthermore, we exclude LSBGs with suspicious HI profiles that indicate potential interactions with other galaxies, serious asymmetric profiles, or profiles with data points affected by poor data quality, as illustrated in Fig. 1. The first two scenarios may suggest an in-equilibrium state of the galaxy, leading to inaccurate estimation of the rotation velocity.
Figure 1: Examples of suspicious HI spectra from Haynes et al. (2018). $V_{\rm h}$ denotes the heliocentric velocity, and $W_{\rm 50}$ denotes the 50% peak width of an HI spectrum, obtained by Haynes et al. (2018). Panel (a): a profile showing potential interactions with other galaxies. Panel (b): a seriously asymmetric profile. Panel (c): a profile containing bad points due to data quality.
Due to the bad spatial resolution of ALFALFA, if a galaxy has a close companion galaxy, its $\rm HI$ spectrum may be contaminated by the companion. To eliminate this possibility, we remove the galaxies with companions within the projected radii of $3.8^{\prime}$ (i.e., approximately the beam size of Arecibo) and radial velocity differences of $\Delta v<500{\rm km/s}$. The redshift of a LSBG is estimated from the central velocity of its $\rm HI$ spectrum  [^1] While the distances of our sample galaxies are obtained directly from Haynes et al., 2018, please refer to their work for more details. , while the redshift of the neighbouring galaxies comes from the SpecObj and Photoz databases of SDSS. Finally, 204 LSBGs are left in our sample.
We also calculate the kurtosis coefficient, $k_{\rm 4}$, of the HI spectrum (Papastergis et al., 2016; El-Badry et al., 2018) for each galaxy, as
@@
Here, $\nu$ and $F(\nu)$ represent the frequency and corresponding flux density, respectively. $\bar{\nu}$ is defined as,
@@
where $\nu_{\rm min}$ and $\nu_{\rm max}$ correspond to the minimal and maximum frequencies between which $W_{20}$ (see Section 3 for the definition of $W_{20}$) is measured, respectively. $k_{\rm 4}$ can distinguish the single-horned and double-horned spectra (see, e.g., Fig. 1 of El-Badry et al., 2018). The galaxies with single-horned HI lines may have kinematics strongly influenced by velocity dispersion (El-Badry et al., 2018) or beam smearing, or HI disks that are not extended enough to trace the flat part of the rotation curve (Papastergis et al., 2016). As a consequence, their HI-widths cannot represent the rotation velocities (McGaugh, 2012; Verheijen, 1997). Therefore, we perform a cut at the HI spectrum, $k_{\rm 4}<-1$, in order to remove the galaxies with single-horned HI lines (see Fig. 3 of El-Badry et al., 2018 for the threshold).
@@
Figure 2: A comparison of properties of selected LSBGs and HSBGs. The green-filled circles represent LSBGs (the dark-green and light-green depict the galaxies with apparent axis ratios $b/a>0.3$ and $b/a\leq 0.3$, respectively), while the grey-filled circles represent HSBGs. Panel (a): axis ratio $b/a$ versus mean surface density $\braket{\Sigma}_{*,\rm{eff}}$. Panel (b): $g-i$ color versus gas fraction $f_{\rm{g}}$. Panel (c): scale length $R_{\rm{d}}$ versus baryonic mass $M_{\rm{b}}$. Panel (d): $v_{\rm rot}$ versus $f_{\rm b}(R_{\rm HI})$. Please refer to Section 2 and Section 5.2 for the definitions of these parameters. The black-solid line and the blue-shaded region in panel d represent the best-fitting linear line and the bin used for performing the two-sample Kuiper tests, respectively, as described in Section 5.2.
@@ 3 Method
The baryonic mass of a galaxy is mainly composed of the stellar mass and gas mass. The stellar mass $M_{\rm*}$ can be estimated with the galaxy luminosity in the $W1$-band (denoted as $L_{W1}$) from the Wide-field Infrared Survey Explorer (Wright et al., 2010), using the method proposed by McGaugh & Schombert (2015), as
@@
As for the gas mass, $M_{\rm g}$, we estimate it by adding up the masses of HI and helium, where $M_{\rm HI}$ is the HI mass (see Haynes et al., 2018 for the method of estimating $M_{\rm HI}$). The baryonic mass of a galaxy is then estimated as $M_{\rm b}=M_{\rm*}+M_{\rm g}$. Note that $M_{\rm b}$ does not include the hot ionized gas and molecular gas, as their fractions are expected to be small compared to the HI fraction (McGaugh, 2012; Zhong et al., 2023), and thus are negligible.
We use the $20\%$ peak width of an HI spectrum (Guo et al., 2020; Hu et al., 2023; Rong et al., 2024, $W_{\rm 20}$;; see the Method section of Guo et al. (2020) for the method of estimation of $W_{\rm 20}$)to calculate the rotation velocity $v_{\rm rot}$. Guo et al. (2020) have demonstrated that $W_{\rm 20}$ can accurately trace galactic $v_{\rm rot}$, which can be estimated as,
@@
where $i$ is the inclination angle. $W_{20}$ employed in equation (7) has been corrected for instrumental broadening and redshift (Guo et al., 2020; Catinella et al., 2012; Kent et al., 2008). The errors of $W_{\rm 20}$ are estimated following the method described by Guo et al. (2020). Please refer to the Method section of their paper for details. In equation (7), we assume an alignment between the inclination angles of the stellar disk and HI disk in a galaxy and then use the inclination angle of the stellar disk to estimate the rotation velocity. The inclination is estimated from the $g-$band apparent axis ratio, as (e.g., Guo et al., 2020; Hu et al., 2023; Du et al., 2019; Rong et al., 2024; Begum et al., 2008), $\sin i=\sqrt{\frac{1-(b/a)^{2}}{1-q_{0}^{2}}}$, where $q_{0}$ denotes the intrinsic axis ratio seen edge-on. We take $q_{0}=0.2$, which is commonly used in the previous studies (e.g., Du et al., 2019; Guo et al., 2020; Giovanelli et al., 1997; Tully et al., 2009). For those edge-on galaxies with $b/a\leq q_{0}$, we set $\sin i=1$.
Previous studies have revealed that the stellar disk and gas disk in galaxies may not be perfectly co-planar, but often exhibit a small inclination difference of ${\rm\delta}i<20^{\circ}$ (e.g., Starkenburg et al., 2019; Guo et al., 2020; Gault et al., 2021). In order to take this misalignment into account, we assume ${\rm\delta}i$ following a Gaussian distribution centred at $0^{\circ}$ with a standard deviation of $\sigma_{i}=20^{\circ}$. We treat $\sigma_{i}=20^{\circ}$ as the uncertainty associated with $i$ in our study. Note that we do not apply corrections for the velocity dispersion in equation (7), since the effects of the dispersion are severe under $v_{\rm rot}\lesssim 40\ {\rm km/s}$ (summarized by Lelli, 2022) while $v_{\rm rot}$ of our sample galaxies are well above $\sim 50\ {\rm km/s}$.
4 BTFR
In Fig. 3, we observe that the distributions of LSBGs and HSBGs exhibit no significant deviation in the $v_{\rm rot}-M_{\rm b}$ diagram. We impose a power-law equation,